Which pair has equally likely outcomes? List the letters of the two choices below which have equal probabilities of success, separated by a comma. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards).
A. rolling a sum of 11 on two fair six sided dice
B. drawing a red nine out of a standard 52 card deck given it’s not a face card or an Ace.
C. rolling a sum of 10 on two fair six sided dice
D. rolling a sum of 7 on two fair six sided dice
E. drawing a six out of a standard 52 card deck given it’s not a face card or an Ace.

Found 2 solutions by ewatrrr, edjones:
Answer by ewatrrr(24785)

 
Hi,
A. rolling a sum of 11 on two fair six sided dice: P = 2/36 = 1/18
B. drawing a red nine out of a standard 52 card deck given it’s not a face card or an Ace.
P(A|B) = P(A and B)/P(B) = 2/36 = 1/18
C. rolling a sum of 10 on two fair six sided dice: P = 3/36 = 1/12
D. rolling a sum of 7 on two fair six sided dice: p = 3*36 = 1/6
E. drawing a six out of a standard 52 card deck given it’s not a face card or an Ace.
P(A|B) = P(A and B)/P(B) = 4/36 = 1/9
___1_2_3__4__5__6
1|_2_3_4__5__6__7
2|_3_4_5__6__7__8
3|_4_5_6__7__8__9
4|_5_6_7__8__9_10
5|_6_7_8__9_10_11
6|_7_8_9_10_11_12

Answer by edjones(8007)

(A, B) 2/36=1/18